OpenStudy (anonymous):

WILL GIVE MEDAL Which of the ordered pairs in the form (x, y) is a solution of this equation? 3x-y/4=11 (3, –8), (4, 4) A. The first is a solution, but the second is not. B. Neither is a solution. C. The first is not a solution, but the second is. D. Both are solutions.

OpenStudy (anonymous):

@jojokiw3

OpenStudy (danjs):

test the points out to see if the equation is true (x,y) = (3, -8) 3*3 - (-8)/4 = 9 + 2 = 11 true that point works, try the second one the same way

OpenStudy (anonymous):

ok what next? @DanJS

OpenStudy (anonymous):

@LoveIt @daisyduck04

OpenStudy (anonymous):

Uh is this middle school math?

OpenStudy (anonymous):

yes @LoveIt

OpenStudy (danjs):

do the same thing i did for the first point

OpenStudy (anonymous):

i did

OpenStudy (danjs):

(x,y) = (4,4), does it work

OpenStudy (anonymous):

no?

OpenStudy (danjs):

3*x - y/4 = 11 for point (x,y)=(4,4), put x=4, and, y=4 into that, see if it remains true or not